# CAPE: Covariate-Adjusted Pre-Training for Epidemic Time Series Forecasting

### v2

@@ 4.4 Transferability
Cross-Location. We include measles data from the USA in the pre-training dataset. To evaluate our model’s ability to adapt to cross-region data, we incorporate measles outbreak data from the UK into the downstream datasets. As shown in Table 4, the pre-trained CAPE outperforms the non-pre-trained version by 17.48%. While we pre-train our model with influenza data from the USA, the zero-shot evaluation on the influenza outbreak in Japan also shows superior performance, underscoring the crucial role of pre-training in enabling generalization to novel regions.
Cross-Disease. While we include various types of diseases in our pre-training dataset, novel diseases including Dengue (non-respiratory) and COVID-19 that are unseen in the pre-training stage are incorporated during the downstream evaluation. The ability of our model to adapt to novel diseases is proven compared to the version not pre-trained on the Dengue dataset (Table 4), improving which by 23.24%, as well as the superior zero-shot performance on the COVID dataset (Table 3), which surpasses the MOMENT that is not pre-trained on other diseases by 72.60%.
Cross-Interval. While we only pre-train using weekly-sampled data, our model outperformed the non-pre-trained version on the irregularly sampled Dengue dataset, demonstrating robustness to different time intervals. Additionally, on the daily-sampled COVID-19 dataset, our model maintained strong zero-shot performance, further illustrating its ability to generalize across varying temporal resolutions.
@@ 4.5 Deeper Analysis
Impact of Pre-Training Epochs. Evaluating four downstream datasets (Figure 3), we find that increasing pre-training epochs consistently improves performance on Measles and COVID datasets but degrades it for ILI USA. Additionally, models with more environment states $K$ perform better as pre-training epochs increase.
Figure 3: Downstream performance with different numbers of environments and pre-training epochs.
Impact of Pre-Training Materials. We examine potential biases in our pre-training dataset by splitting it into respiratory and non-respiratory diseases. As shown in Figure 4, with similar volumes of pre-training data, the model performs better on downstream datasets when their disease types align with the pre-training data (e.g., respiratory diseases). However, the size of the pre-training material has a more significant impact on downstream performance.
Figure 4: Downstream performance variation when the model is pre-trained with either respiratory or non-respiratory diseases only.
Impact of Pre-Training Material Scale. To explore how the pre-training material scale affects downstream performance, we scaled the original pre-training dataset and test on downstream datasets. As shown in Figure 5, a sudden performance boost is observed at around a 60% reduction for both Measles and Dengue datasets.
Figure 5: Downstream performance across pre-training ratios.
Tackling Distribution Shift. In this study, distribution shifts refer to changes in infection patterns observed from the training set to the test set. To evaluate distribution shifts, we compute the Central Moment Discrepancy (CMD) score Zellinger et al. (2017) between training and test distributions for each disease (see Appendix A.8). Figure 6 shows that our model with environment estimation achieves the lowest CMD score, demonstrating its effectiveness in mitigating the impact of temporal distribution shifts.
Figure 6: We report the CMD scores of the embeddings produced by CAPE with and without environment estimation, which quantify distributional differences between the training and test sets.
Disentangling Disease Dynamics. We validate our model’s ability to capture intrinsic disease dynamics by extracting latent embeddings from various datasets and computing the Davies-Bouldin Index (DBI) for each pair. As shown in Figure 7, CAPE consistently achieves lower DBI scores than PatchTST across all pairs, demonstrating its superior effectiveness in distinguishing diseases and separating disease-specific patterns from environmental influences.
Figure 7: Davies-Bouldin Index score between the embeddings of each pair of downstream datasets, output by the pre-trained model without fine-tuning. A visualization is shown in Appendix A.9.
@@ 5 Conclusion
We present Covariate-Adjusted Pre-Training for Epidemic time series forecasting, showcasing the benefits of pre-training and environment modeling. While leveraging pre-training materials, CAPE explicitly learns latent representations of the environment and performs backdoor adjustment. Extensive experiments validate CAPE’s effectiveness in various settings, including few-shot and zero-shot.
@@ Appendix A Appendix
A.1 Theoretical Analysis
A.1.1 Derivation for do-operation
@@ 2. Insertion/Deletion of Actions:
Step 1: Action/Observation Exchange. Using $(\hat{Y}\perp\!\!\!\perp X|Z)$ in $\mathcal{A}^{X}$, we apply the exchange rule:
@@
Step 2: Insertion/Deletion of Actions. Using $(Z\perp\!\!\!\perp X)$ in $\mathcal{A}^{X}$, we simplify:
@@
A.1.2 Derivation for Variational Lower Bound
@@
$=\sum_{z}q_{\phi}(Z=z|X)\log p_{\Theta}(\hat{Y}|X,Z=z)$
$\quad\quad\quad\quad\quad\quad\quad-\sum_{z}q_{\phi}(Z=z|X)\log\frac{q_{\phi}(Z=z|X)}{p(Z=z)}$
@@
$\log\left(\frac{\pi_{i}^{t}}{\pi_{i}^{t+1}}\right)$ $=\log\left(1+\frac{\pi_{i}^{t}-\pi_{i}^{t+1}}{\pi_{i}^{t}}\right)$ (17)
$=\log\left(1-\frac{\Delta\pi_{i}}{\pi_{i}^{t}}\right)\approx-\frac{\Delta\pi_{i}}{\pi_{i}^{t}}-\frac{1}{2}\left(\frac{\Delta\pi_{i}}{\pi_{i}^{t}}\right)^{2}.$
@@
$\text{KL}\left(q_{\phi}^{t}\parallel q_{\phi}^{t+1}\right)$ $\approx\sum_{i}\pi_{i}^{t}\left(-\frac{\Delta\pi_{i}}{\pi_{i}^{t}}-\frac{1}{2}\left(\frac{\Delta\pi_{i}}{\pi_{i}^{t}}\right)^{2}\right)$ (18)
$=-\sum_{i}\Delta\pi_{i}-\frac{1}{2}\sum_{i}\frac{(\Delta\pi_{i})^{2}}{\pi_{i}^{t}}.$
@@
$\hskip 142.26378pt\cdot\frac{q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})}{q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})}$
$\geq\sum_{\mathbf{x}}p(X=\mathbf{x})\sum_{\mathbf{z}}q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})$
$\hskip 42.67912pt\cdot\log{p_{\Theta}(\hat{Y}|X=\mathbf{x},Z=z)}\frac{p(Z=z)}{q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})}$
@@
$=\sum_{\mathbf{x}}p(X=\mathbf{x})\sum_{\mathbf{z}}q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})$
$\hskip 119.50148pt\cdot\log{p_{\Theta}(\hat{Y}|X=\mathbf{x},Z=z)}$
$\hskip 58.32823pt+\sum_{\mathbf{x}}p(X=\mathbf{x})\sum_{\mathbf{z}}q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})$
$\hskip 123.76965pt\cdot\log{\frac{p(Z=z)}{q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})}}$
@@
$=\sum_{\mathbf{x}}p(X=\mathbf{x})\sum_{z}q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})$
$\hskip 136.5733pt\cdot\log{\frac{p(Z=z)}{q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})}}$
@@
$\hskip 69.70915pt-\sum_{\mathbf{x}}p(X=\mathbf{x})\sum_{z}q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})$
$\hskip 136.5733pt\cdot\log{q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})}$
$=\sum_{z}\log{p(Z=z)}\sum_{\mathbf{x}}p(X=\mathbf{x})q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})$
$\hskip 176.407pt-H_{\phi_{t}}(Z|X)$
@@
$E_{p(X)}\log{p_{\Theta}(\hat{Y}|do(X))}$ (24)
$\geq\mathbb{E}_{p(X)}\mathbb{E}_{q^{t+1}_{\phi}(Z|X)}\log{p_{\Theta}(\hat{Y}|X,Z)}-C$
Therefore we only focus on maximizing the first term, which becomes minimizing the reconstruction loss $\mathcal{L}_{recon}$. In practice, we make $q^{t+1}_{\phi}$ active during this step while posing an L2 norm on its parameters to better optimize the environment estimator as well as switching to a different environment distribution space. During training, as proven in Proof A.1.3, the updated $q^{t+1}_{\phi_{t+1}}$ eventually approximates the $q^{t+1}_{\phi_{t}}$.
A.2 Pre-training and Downstream Datasets Details
@@
In addition, we collect seven datasets of different types of diseases from diverse sources for downstream evaluations, which are all normalized without further processing. A summary of the downstream datasets is shown in Figure 6.
@@
A.3 Implementation Details
Settings. We adopt an input length of 36 Wu et al. (2023); Wang et al. (2024b) and a patch size of 4 for applicable models. For the environment estimator defined in Eq. (5), a shared weight $w_{k}$ is used for all environment representations. All results are evaluated using Mean Squared Error (MSE).

### v3

@@ 4.4 Transferability
Cross-Location: We include measles data from the USA in the pre-training dataset. To evaluate our model’s ability to adapt to cross-region data, we incorporate measles outbreak data from the UK into the downstream datasets. As shown in Table 4, the pre-trained CAPE outperforms the non-pre-trained version by 17.48%. While we pre-train our model with influenza data from the USA, the zero-shot evaluation on the influenza outbreak in Japan also shows superior performance, underscoring the crucial role of pre-training in enabling generalization to novel regions. Cross-Disease: While we include various types of diseases in our pre-training dataset, novel diseases including Dengue (non-respiratory) and COVID-19 that are unseen in the pre-training stage are incorporated during the downstream evaluation. The ability of our model to adapt to novel diseases is proven compared to the version not pre-trained on the Dengue dataset (Table 4), improving which by 23.24%, as well as the superior zero-shot performance on the COVID dataset (Table 4), which surpasses the MOMENT that is not pre-trained on other diseases by 72.60%. Cross-Interval: While we only pre-train using weekly-sampled data, our model outperformed the non-pre-trained version on the irregularly sampled Dengue dataset and on the daily-sampled COVID-19 dataset, demonstrating robustness to different time intervals.
@@ 4.5 Deeper Analysis
Impact of Pre-Training Epochs: Evaluating four downstream datasets (Figure 5), we find that increasing pre-training epochs consistently improves performance on Measles and COVID datasets but degrades it for ILI USA. Additionally, models with more environment states $K$ perform better as pre-training epochs increase. Impact of Pre-Training Materials: We examine potential biases in our pre-training dataset by splitting it into respiratory and non-respiratory diseases. As shown in Figure 4, with similar volumes of pre-train data, the model performs better when the tested disease types align with the pre-train data. However, the size of the pre-train material has a stronger impact. Impact of Pre-Training Material Scale: To explore how the pre-training material scale affects downstream performance, we scaled the original pre-training dataset and test on downstream datasets. As shown in Figure 6, a sudden performance boost is observed at around a 60% reduction for both Measles and Dengue datasets.
Figure 4: Downstream performance when the model is pre-trained with either respiratory or non-respiratory data.
Tackling Distribution Shift: We define distribution shifts as changes in infection patterns from training to test data. We compute the Central Moment Discrepancy (CMD)(Zellinger et al., 2017) between training and test distributions for each disease. As shown in Figure 7, CAPE with environment estimation achieves the lowest CMD scores, highlighting its effectiveness in mitigating distribution shifts. Disentangling Disease Dynamics: We assess CAPE’s ability to capture intrinsic disease dynamics by extracting latent embeddings and computing the Davies-Bouldin Index (DBI) across different dataset pairs. As shown in A.13, CAPE consistently yields lower DBI scores than PatchTST, indicating better separation of disease-specific patterns from environmental influences.
Figure 5: Downstream performance vs. environments distribution. More results in A.10.
Figure 6: Downstream performance across pre-training ratios. More datasets are in  A.11.
Figure 7: CMD scores w/ and w/o environment estimate between train/test sets. (A.12)
@@ 5 Conclusion
We present CAPE, a covariate-adjusted pre-training framework for epidemic forecasting that captures environment-invariant disease dynamics by combining pre-training with implicit environment modeling grounded in causal analysis. By learning latent environments and applying backdoor adjustment, CAPE reduces distribution shifts. Extensive experiments across diseases, regions, and scenarios demonstrate strong generalization, with consistent gains in few-shot and zero-shot settings. Future directions can include incorporating uncertainty measurement, scaling to larger datasets, and providing interpretable environment representations.
@@ Appendix A Appendix
A.1 Potential Impact
This paper advances the interdisciplinary fields of machine learning and epidemiology by enhancing the accuracy of epidemic forecasting in data-limited settings. We outline the potential impacts as follows:
Early Insights. We provide novel insights into how pre-training and environment modeling improve epidemic forecasting. Our results demonstrate that pre-training significantly enhances model accuracy, with gains increasing as more pre-training data is incorporated. This finding paves the way for future research to develop foundational models in epidemic analysis using larger datasets. Additionally, we confirm the importance of accounting for both inherent disease dynamics and environmental factors to achieve robust forecasting performance.
Social Impact. Epidemic time series data are often sparse due to limited sampling rates, hindering public health organizations’ ability to accurately predict infections during novel disease outbreaks. This paper addresses this challenge by showcasing the few-shot and zero-shot forecasting capabilities of pre-trained models. These capabilities can provide powerful tools for early warning and timely intervention, ultimately supporting more effective public health responses and safeguarding communities against emerging infectious diseases.
A.2 Limitations
While CAPE shows strong performance, it has several limitations. First, it does not currently incorporate uncertainty estimation, which is important for risk-aware decision-making. Second, its effectiveness may be constrained by the scale and diversity of available data. Lastly, the learned environment representations lack interpretability, limiting their transparency and potential for insight in public health contexts.
Though not explicitly designed for uncertainty, CAPE supports Dropout Injection (Ledda et al., 2023) at inference. We use 25 stochastic forward passes on the ILI USA dataset to compute 95% confidence intervals:
Table 5: CAPE Forecasting Results with 95% Confidence Intervals
Horizon CAPE 95% CI
1 $0.174\pm 0.0036$
2 $0.235\pm 0.0016$
4 $0.293\pm 0.0012$
8 $0.433\pm 0.0010$
A.3 Theoretical Analysis
A.3.1 Derivation for do-operation
@@ 2. Insertion/Deletion of Actions:
Step 1: Action/Observation Exchange. Using $(\hat{Y}\perp\!\!\!\perp X|Z)$ in $\mathcal{A}_{X}$, we apply the exchange rule:
@@
Step 2: Insertion/Deletion of Actions. Using $(Z\perp\!\!\!\perp X)$ in $\mathcal{A}_{X}$, we simplify:
@@
A.3.2 Derivation for Variational Lower Bound
@@
$=\sum_{z}q_{\phi}(Z=z|X)\log p_{\Theta}(\hat{Y}|X,Z=z)-\sum_{z}q_{\phi}(Z=z|X)\log\frac{q_{\phi}(Z=z|X)}{p(Z=z)}$
@@
$\log\left(\frac{\pi_{i}^{t}}{\pi_{i}^{t+1}}\right)$ $=\log\left(1+\frac{\pi_{i}^{t}-\pi_{i}^{t+1}}{\pi_{i}^{t}}\right)=\log\left(1-\frac{\Delta\pi_{i}}{\pi_{i}^{t}}\right)\approx-\frac{\Delta\pi_{i}}{\pi_{i}^{t}}-\frac{1}{2}\left(\frac{\Delta\pi_{i}}{\pi_{i}^{t}}\right)^{2}.$ (13)
@@
$\text{KL}\left(q_{\phi}^{t}\parallel q_{\phi}^{t+1}\right)$ $\approx\sum_{i}\pi_{i}^{t}\left(-\frac{\Delta\pi_{i}}{\pi_{i}^{t}}-\frac{1}{2}\left(\frac{\Delta\pi_{i}}{\pi_{i}^{t}}\right)^{2}\right)=-\sum_{i}\Delta\pi_{i}-\frac{1}{2}\sum_{i}\frac{(\Delta\pi_{i})^{2}}{\pi_{i}^{t}}.$ (14)
@@
$=\sum_{\mathbf{x}}p(X=\mathbf{x})\sum_{\mathbf{z}}\log{p_{\Theta}(\hat{Y}|X=\mathbf{x},Z=z)}p(Z=z)\cdot\frac{q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})}{q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})}$
$\geq\sum_{\mathbf{x}}p(X=\mathbf{x})\sum_{\mathbf{z}}q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})\cdot\log{p_{\Theta}(\hat{Y}|X=\mathbf{x},Z=z)}\frac{p(Z=z)}{q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})}$
@@
$=\sum_{\mathbf{x}}p(X=\mathbf{x})\sum_{\mathbf{z}}q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})\cdot\log{p_{\Theta}(\hat{Y}|X=\mathbf{x},Z=z)}$
$\hskip 58.32823pt+\sum_{\mathbf{x}}p(X=\mathbf{x})\sum_{\mathbf{z}}q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})\cdot\log{\frac{p(Z=z)}{q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})}}$
@@
$=\sum_{\mathbf{x}}p(X=\mathbf{x})\sum_{z}q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})\cdot\log{\frac{p(Z=z)}{q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})}}$
@@
$\hskip 28.45274pt-\sum_{\mathbf{x}}p(X=\mathbf{x})\sum_{z}q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})\cdot\log{q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})}$
$=\sum_{z}\log{p(Z=z)}\sum_{\mathbf{x}}p(X=\mathbf{x})q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})+H_{\phi_{t}}(Z|X)$
@@
$E_{p(X)}\log{p_{\Theta}(\hat{Y}|do(X))}\geq\mathbb{E}_{p(X)}\mathbb{E}_{q^{t+1}_{\phi}(Z|X)}\log{p_{\Theta}(\hat{Y}|X,Z)}-C$ (20)
Therefore we only focus on maximizing the first term, which becomes minimizing the reconstruction loss $\mathcal{L}_{recon}$. In practice, we make $q^{t+1}_{\phi}$ active during this step while posing an L2 norm on its parameters to better optimize the environment estimator as well as switching to a different environment distribution space. During training, as proven in Proof A.3.3, the updated $q^{t+1}_{\phi_{t+1}}$ eventually approximates the $q^{t+1}_{\phi_{t}}$.
A.4 Pre-train and Downstream Datasets Details
Due to incomplete records and limited non-U.S. data, we curated diverse, high-quality datasets emphasizing:
  1. Epidemic variation: Decades of weekly data (visualization) reflect multi-wave trends and interventions.
  2. Geographic spread: Covers 50 U.S. states and 25 non-U.S. Dengue regions.
  3. Climate diversity: Implicit variation (e.g., tropical Florida vs. temperate Maine).
  4. Data quality: Manually curated, gap-free datasets (from CDC, ProjectTycho, etc.).
We apply z-score normalization for all datasets to align disease data with zero-mean.
A.4.1 Pre-train Datasets
@@
A.4.2 Downstream Datasets
In addition, we collect seven datasets of different types of diseases from diverse sources for downstream evaluations, which are all normalized without further processing. A summary of the downstream datasets is shown in Table 7.
@@ A.4.2 Downstream Datasets
A more detailed description of each dataset is shown below:
  1. ILI USA  (Centers for Disease Control and Prevention, 2023a): The weekly influenza-like-illness infection was reported by the CDC in the United States. We use the national-level infection counts from 2002 to 2020, which include various disease such as H1N1, H3N2v, etc.
  2. ILI Japan  (National Institute of Infectious Diseases, 2023): This dataset is collected from the Infectious Diseases Weekly Report (IDWR) in Japan, which contains national counts of weekly influenza-like-illness infections from August 2012 to March 2019.
  3. Measles  (Lau et al., 2020a): The measles dataset contains biweekly measles infections in England from 1906 to 1948.
  4. Dengue  (OpenDengue, 2023): OpenDenue aims to build and maintain a database of dengue case counts for every dengue-affected country worldwide since 1990 or earlier. We selected 23 countries for the experiment, which reports daily to weekly infections.
  5. RSV  (Centers for Disease Control and Prevention, 2023c): The Respiratory Syncytial Virus (RSV) infections in the US are reported by the RSV-NET from CDC. We use the weekly infections across 13 states from 2016 to 2024.
  6. MPox  (Centers for Disease Control and Prevention, 2023b): The clade II MPox case trends data in the US is reported by CDC. We use the nationwide weekly infections from 2022 to 2024.
  7. COVID  (Dong et al., 2020): The original data is from the Novel Coronavirus Visual Dashboard operated by the Johns Hopkins University Center for Systems Science and Engineering (JHU CSSE). We use the daily COVID-19 infections collected by JHU from 2020 to 2022 across 16 states.
A.5 Implementation Details
Motivation for Our Settings. Our primary focus is on pre-training epidemic forecasting models using temporal (time series) data rather than spatiotemporal data. This design choice is motivated by the following considerations:
  1. We aim to establish the foundation for epidemic pre-training in the temporal setting, which remains unexplored — only one prior work  (Kamarthi & Prakash, 2023) addresses this area, and it also focuses on temporal setting. As shown in Sections 4.2-4.5, we address critical questions around generalization, few-shot/zero-shot performance, and pre-training dynamics — topics that remain open even without spatial context.
  2. Temporal models are broadly applicable and more data-efficient, especially when spatial data is unavailable or unreliable. Many real-world epidemic datasets lack well-defined spatial graphs, and building them (e.g., from mobility or administrative data) is costly and complex, particularly at scale. These inconsistencies also hinder fair comparisons between temporal and spatiotemporal models.
  3. Our framework is extensible to spatiotemporal modeling. Specifically, the temporal input can be replaced with graph structured data, and the predictor can incorporate graph-based encoders. Exploring this direction is exciting future work, but we believe temporal pre-training is a crucial first step toward that goal.
Zero-Shot. Once pre-trained, our CAPE framework can be directly utilized for zero-shot forecasting where the model remains frozen and no parameter is updated. Similar to the MOMENT model (Goswami et al., 2024), we retain the pre-trained reconstruction head and mask the last patch of the input to perform forecasting: $\hat{\mathbf{y}}=\hat{\mathbf{x}}_{[T-c:T]}$.
Model Details. We design our model by stacking 4 layers of the CAPE encoder, each with a hidden size of 512 and 4 attention heads. For environment representations, we incorporate 16 distinct environments, each encoded with a size of 512. To ensure a fair comparison, PatchTST is configured with the same number of layers and hidden size as our CAPE-based model. For all other baseline models, we adopt the architectures as reported in previous studies (Wang et al., 2024a; Kamarthi & Prakash, 2023; Panagopoulos et al., 2021).
Training Details:
  1. We adopt an input length of 36 (Wu et al., 2023; Wang et al., 2024b) and a patch size of 4 for applicable models. For the environment estimator defined in Eq. (4), a shared weight $w_{k}$ is used for all environment representations. All results are evaluated using Mean Squared Error (MSE).
  2. CAPE follows the general EM framework, whose convergence is well-studied (Obata et al., 2024; Qu et al., 2019; Wu, 1983). To ensure stable convergence, we use a small learning rate (1e-5) with L2 regularization, train for 150 epochs, and select the model with the lowest validation error.
  3. For the training process, we pre-train CAPE, PEM, and PatchTST on a single Nvidia A100 GPU. During pre-training, we utilize only 70% of the available training data, specifically the first 70% of the dataset for each disease category. We set the learning rate to $1\times 10^{-5}$ and apply a dropout rate of 0.1 to prevent overfitting. In the CAPE pre-training strategy, we assign a weight of 0.5 to $\alpha$ to balance the contributions of contrast loss to the whole loss function. A detailed illustration of our pre-training strategy is shown in Algorithm 1.
  4. After pre-training, we fine-tune the entire model using a single Nvidia K80 GPU, maintaining the same hyperparameter settings for consistency. The best-performing model is selected based on its performance on the validation set. Similarly, for all baseline models, we train each until convergence and select the optimal model based on validation set performance for the subsequent test.